OpenStudy (anonymous):

Prove , cos20cos40cos80=1/8 . Angles are in degrees .

OpenStudy (anonymous):

by talk about "trickonometry" who thought up this question?

OpenStudy (anonymous):

one gimmick is to call \[x=\cos(20)\cos(40)\cos(80)\] then write \[\sin(20)x=\sin(20)\cos(20)\cos(40)\cos(80)\]

OpenStudy (anonymous):

now use the "double angle" formula for sine twice. that is use \[\sin(2x)=2\sin(x)\cos(x)\] so you get \[\sin(20)x=\frac{1}{2}\sin(40)\cos(40)\cos(80)\] then again and get \[\sin(20)x=\frac{1}{4}\sin(80)\cos(80)\] a nd and again to get \[\sin(20)x=\frac{1}{8}\sin(160)\] finally a miracle occurs, namely that \[\sin(20)=\sin(160)\] since 20+160=180 giving you \[\sin(20)x=\frac{1}{8}\sin(20)\] cancel to get \[x=\frac{1}{8}\]

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

But what about sin10sin50sin70 ?

OpenStudy (anonymous):

Figured it out , thank you again !